OpenStudy (anonymous):

Using elimination what is the solution of the system? -5x - 2y = 4 3x + 6y = 6

OpenStudy (anonymous):

you could start by eliminating the y's by multiplying the 1st equation by 3. Yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

-15x 12y=24 but what sing goes between -15x and 12y?

OpenStudy (anonymous):

-, look at it this way 3(-5x-4y=4) is the same as 3(-5x)+3(-4y)=3(4) 3(4) is not 24 by the way

OpenStudy (anonymous):

the answer they give are a) (-5,10.5) b) (-3,2.5) c) (-1.5,1.75) d) (1.5,0.25)

OpenStudy (anonymous):

@emma151 rewite equation one so that you can subtract it from equation two by multiplying by 3.. waht do you get?

OpenStudy (anonymous):

so multiply first equation by 3?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

-15 - 6 = 12?

OpenStudy (anonymous):

yes but with the letters so -15x-6y=12

OpenStudy (anonymous):

so now we have these equations: -15x-6y=12 3x+6y=6 which term can you eliminate? the y or the x? and how?

OpenStudy (anonymous):

@emma151

OpenStudy (anonymous):

The 6y because they are the same now

OpenStudy (anonymous):

yes, good.so you add the 1st equation to the second

OpenStudy (anonymous):

write down what you get :)

OpenStudy (anonymous):

-12x = 18

OpenStudy (anonymous):

x=1.5 right

OpenStudy (anonymous):

yes, now plug it into any equation and you get your y :)

OpenStudy (anonymous):

by the way x= -1,5 since its 18/-12

OpenStudy (anonymous):

@emma151

OpenStudy (anonymous):

y=1.75

OpenStudy (anonymous):

Thank you @Servet! Thnak you @bronzegoddess!

OpenStudy (anonymous):

you are welcome :)